# Incremental Joint Learning of Depth, Pose and Implicit Scene Representation on Monocular Camera in Large-scale Scenes

### v2

@@ Abstract
Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, autonomous vehicles. However, most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input. Accurate depth input is impossible to get in real-world large-scale scenes. (b) inaccurate pose estimation. Existing methods are not robust enough with the growth of cumulative errors in large scenes and long sequence. (c) insufficient scene representation capability. A single global radiance field lacks the capacity to effectively scale to large-scale scenes. To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale scene reconstruction. A vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation. For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes. In terms of implicit scene representation, we propose an incremental scene representation method to construct the entire large-scale scene as multiple local radiance fields to enhance the scalability of 3D scene representation. Extended experiments have been conducted to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction. The code will be open-sourced upon paper acceptance.
@@
Neural Radiance Fields, Incremental Joint Learning, Pose Estimation, Large-scale Scenes.
@@ I Introduction
In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics, autonomous driving. Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [1]. The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [2], which makes it challenging to use in real-world deployments due to the long processing time. Structure from motion (SFM) is also not robust and accurate in hand-held scenes. Some recent works are proposed to simultaneously estimate camera poses and neural implicit representation, such as BARF [3], SC-NeRF [4], and so on [5, 6, 7]. However, existing methods come with the assumptions of a fixed-size map, which renders them useless in the real-world application.
Figure 1: Qualitative results on the Tanks and Template datasets [8] on Ballroom sequence. Our method achieve better view synthesis performance than existing methods.
@@
Recent works, such as NoPe-NeRF [9] and LocalRF [10], have been proposed to improve the scene representation, which is closely related to our work. They design a joint learning approach for pose and scene representation. NoPe-NeRF also estimates the scale correction parameters during training. LocalRF progressively estimates the poses of input frames. However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts. Their end-to-end pose estimation are also sensitive to pixel-wise noise variation in longer trajectory sequences.
@@
This paper focuses on large-scale 3D reconstruction, highlighting three key challenges: (a) inaccurate depth scale estimation, crucial for accurate geometry in neural scene representation, particularly challenging in large-scale scenes; (b) inaccurate pose estimation, which current methods struggle with due to error accumulation in large-scale scenes; (c) limited scene representation capability, where existing approaches using a single global model hinder scalability to larger scenes and longer sequences.
@@
To this end, we propose an incremental joint learning method for depth, pose estimation, and scene implicit representation. These three elements are closely interconnected: depth information is crucial for geometric representation of the scene. It can guide the neural point sampling and enhance scene geometry consistency. Accurate pose estimation is also vital for scene representation. So, we design a joint learning framework for these three elements. The depth estimation network adopts a vision transformer-based network to encode input images into high-level features. The transformer backbone utilizes a global receptive field to process features at every stage, which provides a more precise estimation. For the pose estimation network, a feature-metric bundle adjustment (FBA) method is proposed for robust pose estimation. Compared to the existing BA methods, our FBA method effectively avoids the sensitivity of pixel-wise noise variation, enabling robust localization in large-scale scenes. For scene implicit representation, existing methods usually use an MLP to encode the scene, leading to poor scene representation capability. We propose an incremental scene representation method. We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation. The incremental scene representation method can significantly enhance the scene representation capability in large-scale scenes.
@@
  • A novel incremental joint learning method is proposed for accurate depth estimation, pose estimation, and large-scale scene reconstruction.
  • We design an FBA method for the pose estimation module. Our method extracts multi-level features with pixel-wise confidences and incorporates a coarse-to-fine BA strategy to achieve robust and accurate pose estimation in large-scale scenes.
  • An incremental scene representation method is designed, which dynamically initiates local radiance fields trained with frames within a local window. This enables accurate scene reconstruction in arbitrarily long video and large-scale scenes.
@@ II Related work
Depth estimation, pose estimation, and scene representation based on monocular images have been a major problem in robotics and computer vision. Here, we summarize some relevant works.
@@ II-A Joint Learning of Depth and Pose
Some early works [11, 12] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images. [13] use synthetic images to help the disparity learning. Recently, some works have focused on self-supervised learning to jointly estimate pose and depth.  [14, 15, 16, 17, 18, 19] and [20] jointly optimize pose and depth via photometric reprojection error. [21] focus on the scale ambiguity for monocular depth estimation with normalization. [22] proposes a masked GAN network for unsupervised monocular depth and ego-motion estimation. [23] jointly perform depth and camera pose estimation and semantic scene segmentation. However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency. [24] incorporate channelwise attention mechanism spatialwise attention mechanism for accurate depth estimation.
@@ II-B Implicit Scene Representation
Early Novel View Synthesis methods applied interpolations between pixels [25], later works use meshes, Multi-Plane Images, and light field [26]. With the proposal of NeRF [1], many researchers explore combining this implicit method into various applications, such as autonomous driving [27] and robot navigation. [28] integrates neural radiance fields with navigation networks end-to-end as a memory structure for the first time, highlighting the potential applications of neural radiance fields in downstream control domain tasks. [29] design a waypoint-based task decomposition strategy which effectively enhances the reliability of learning-based navigation. The key idea of NeRF is to model the radiance field of a scene with a multi-layer perceptron (MLP). However, the representation of volume densities can not commit the geometric consistency, leading to poor surface prediction for reconstruction tasks. UNISURF [30] and NeuS [31] are proposed to combine world-centric 3D geometry representation with neural radiance fields. They replace the volume density with Signed Distance Field (SDF) values, which can further improve the surface reconstruction. Other methods [32, 33] use various scene geometry representation methods, such as truncated signed distance function, voxel grid.
@@ II-C NeRF with Pose Optimization
The original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments. They require pre-computed camera parameters obtained from SFM and SLAM algorithms [34, 35, 36]. Recently, many researchers have focused on implicit scene representation without pose prior. iNeRF [37] is the first work focused on pose estimation and scene representation. They estimate the camera pose for novel view images with a pre-trained NeRF model. NeRFmm [38] jointly optimised as learnable parameters with NeRF training, through a photometric reconstruction. BARF [3] proposes a coarse-to-fine Bundle Adjustment method to estimate camera poses with neural 3D representation. Some other methods [39, 40, 41, 42] combine semantic feature with NeRF to improve the scene representation and camera tracking. NoPe-NeRF [9] and LocalRF [10] are the most relevant work to ours. NoPe-NeRF learns scale and shift parameters during training and incorporates undistorted monocular depth prior. However, they often fail in large-scale scene representation due to the limited scene representation capacity. Their pose estimation module are not robust and accurate in large-scale scenes, resulting in increased pose drifts.
@@ III Joint learning of depth, pose, and implicit scene representation
In this work, we propose a joint learning framework for depth, pose estimation, and implicit scene representation. The architecture of our framework is shown in Fig. 2. The system only takes in RGB images. There are three networks in our system: depth network, pose network, and scene representation network. We will introduce our three networks separately.
@@ III-A Depth Estimation
In our depth network, we choose the vision transformer network as our backbone. Vision transformer network [43] has been proven to have strong feature learning capabilities. Most large vision models, such as Segment Anything Model [44], choose it as their backbone. Inspired by DPT [45], we leverage vision transformers in place of convolutional neural networks for dense depth prediction. We maintain an encoder-decoder structure for our depth network. For the input images $I\in H\times W\times 3$, these images are decomposed into a sequence of 2D patches, where $(H,W)$ is the height and width of the original image, $(P,P)$ is the resolution of each image patches. We use ResNet50 with group norm and weight standardization to encode these image patches. Our vision transformer module transforms these patches using sequential blocks of multi-headed self-attention (MSA).
@@ III-B Feature-metric BA Pose Estimation
In terms of pose estimation, we propose a FBA network. We show the architecture of our network in Fig. 2. Inspired by [46], each image is encoded into high-level features and uses FBA to get an accurate and robust pose estimation. For every consecutive frames $I_{a},I_{b}$, we use a Unet-like encoder-decoder network to extract $D_{l}$ dimensional feature map $\mathbf{F}^{l}\in\mathbb{R}^{W_{l}\times H_{l}\times D_{l}}$ at each level $l\in\{L,\ldots,1\}$. Those multi-channel image features with decreasing resolution can encode richer semantic information and larger spatial information of the image. Our network can also simultaneously estimate the confidence of each pixel to select the robust features for pose estimation. Compared with the original BA method, this learned representation is robust to large illumination or viewpoint changes, which can provide accurate pose estimation in indoor and outdoor environments. Despite poor initial pose estimates, it can also provide meaningful gradients for successful estimation. Those joint learning methods with classical direct alignment [9, 10] operate on the original RGB image, which is not robust to large-scale scenes and long-term changes, and resorts to Gaussian image pyramids, which still largely limits the convergence to frame-to-frame tracking.
@@
where $\rho(\cdot)$ is a robust cost function with derivative $\rho^{\prime}$, and $\omega_{a}^{i}$ is a per-residual weight. We use the Levenberg-Marquardt (LM) algorithm to iteratively optimize this nonlinear least-squares cost from an initial estimate. Follow [46], each pose update $\delta\in R$ is parametrized on the SE(3) manifold using its Lie algebra.
@@
where $\lambda$ is the damping factor and H is the Hessian matrices. Then, the final pose can be defined by left-multiplication on the manifold.
@@
2:  $p\leftarrow 1$; /$\ast$/ ∗ First camera pose index/$\ast$/∗ /
@@
7:     while $m=0;m<100$ do
8:        Optimize $(\theta_{j})$; /$\ast$/ ∗ Optimize the first local radiance field with the first frame/$\ast$/∗ /
9:        $m\leftarrow m+1$;
10:     end
11:     while $\{R,t\}_{j}<Bound_{j}$ do
12:        $q\leftarrow q+1$;
13:        $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; /$\ast$/ ∗ Append a pose at the end of the trajectory $\ast$/∗ /
@@
15:     end
16:     if $q<P$ then
17:        $j\leftarrow j+1$; /$\ast$/ ∗ A new local radiance field index/$\ast$/∗ /
18:        $\theta_{j}\leftarrow InitializeLR()$ /$\ast$/ ∗ Initialize a new local radiance field/$\ast$/∗ /
19:        $t_{j}\leftarrow t_{q}$; /$\ast$/ ∗ Centered around the last pose/$\ast$/∗ /
20:        $p\leftarrow p+n$; /$\ast$/ ∗ Stop considering the frames in last local radiance field/$\ast$/∗ /
21:     end if
22:     Repeat until all frames are registered.
23:  end
@@
where $\gamma$ is the Huber cost. With coarse-to-fine pose estimation, low-resolution feature maps are responsible for the robustness of the pose estimation, while finer features enhance its accuracy. This feature-metric BA loss weights the supervision of the rotation and translation adaptively for each training example. It greatly enhances the accuracy and robustness of our pose network in large-scale scenes.
@@
Scenes BARF [3] NoPe-NeRF [9] LocalRF [10] Ours
[…]
Tanks and Temples [8] Ballroom 20.66 0.50 0.30 24.97 0.71 0.20 25.29 0.89 0.08 26.50 0.85 0.08
Barn 25.28 0.64 0.24 25.77 0.68 0.22 27.89 0.88 0.11 28.59 0.91 0.11
Church 23.17 0.62 0.27 23.60 0.67 0.23 28.07 0.91 0.07 28.82 0.92 0.08
Family 23.04 0.61 0.28 23.77 0.68 0.22 28.60 0.92 0.007 30.48 0.94 0.05
Francis 25.85 0.69 0.28 29.48 0.80 0.18 31.71 0.93 0.11 33.38 0.93 0.10
Horse 24.09 0.72 0.21 27.30 0.83 0.13 29.98 0.94 0.07 30.07 0.93 0.07
Ignatius 21.78 0.47 0.30 23.77 0.61 0.23 26.54 0.86 0.09 27.33 0.87 0.09
Museum 23.58 0.61 0.28 26.39 0.76 0.17 26.35 0.86 0.16 29.43 0.92 0.12
Mean 23.42 0.61 0.27 26.34 0.74 0.19 27.87 0.91 0.09 29.04 0.90 0.09

### v3

@@ Abstract
Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, and robotics navigation. Existing dense reconstruction methods are primarily designed for small room scenarios, but in practice, the scenes encountered by robots are typically large-scale environments. Most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input. Depth information is crucial for both scene geometry reconstruction and pose estimation. Accurate depth input is impossible to get in real-world large-scale scenes. (b) inaccurate pose estimation. Existing methods are not robust enough with the growth of cumulative errors in large scenes and long sequences. (c) insufficient scene representation capability. A single global radiance field lacks the capacity to scale effectively to large-scale scenes. To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale dense scene reconstruction. For depth estimation, a vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation. For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes and eliminates pose drift. In terms of implicit scene representation, we propose an incremental scene representation method to construct the entire large-scale scene as multiple local radiance fields to enhance the scalability of 3D scene representation. In local radiance fields, we propose a tri-plane based scene representation method to further improve the accuracy and efficiency of scene reconstruction. We conduct extensive experiments on various datasets, including our own collected data, to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction. The code has been open-sourced on https://github.com/dtc111111/incre-dpsr.
Note to Practitioners—In practical robot deployment scenarios, dense scene reconstruction can play a crucial role, such as in precise localization and obstacle avoidance. However, existing scene reconstruction methods are mainly suitable for small-room environments and struggle with larger-scale settings like long corridors or extended sequences. Our method introduces an incremental joint learning framework that simultaneously performs depth estimation, pose estimation, and scene reconstruction. Experimental results demonstrate the accuracy and effectiveness of this approach across different scene types.
@@
3D Scene Reconstruction, Incremental Joint Learning, Neural Radiance Fields, Pose Estimation, Large-scale Scenes.
Figure 1: Qualitative results on the Tanks and Template datasets on Ballroom sequence. We present the visualization results of different methods’ reconstruction results in the four corners of the image. Our method achieve better view synthesis performance than existing methods.
@@ I Introduction
In the past two decades, the demand for scene reconstruction has increased across various applications, including mobile robotics, autonomous driving. Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [1]. However, the original NeRF method is only applicable to small room scenes and cannot perform accurate reconstruction in large-scale scenes, which limits the applicability of these methods in real-world robotic scenarios. In addition, the original NeRF method needs the accurate pose input, which makes it challenging to use in real-world deployments due to the inability to obtain accurate poses in practical robot operating environments. Pre-processing methods, such as Structure from motion (SFM), is also not robust and accurate in large-scale scenes.
@@
Some recent works are proposed to simultaneously estimate camera poses and neural implicit representation, such as BARF [2], SC-NeRF [3], and so on [4, 5, 6]. Recent works, such as NoPe-NeRF [7] and LocalRF [8], have been proposed to further improve the scene representation and pose estimation. NoPe-NeRF estimates the scale correction parameters during training to improve scene representation. LocalRF progressively estimates the poses of input frames to improve the scalability. However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts. Their RGB loss-based pose estimation are also sensitive to pixel-wise noise variation in long trajectory sequences.
@@
In this paper, we focus on large-scale 3D reconstruction without pose information, utilizing only monocular camera input. We highlight three key challenges in large-scale scenes: (a) inaccurate depth scale estimation Depth information is crucial for accurate geometry in neural scene representation, particularly in large-scale scenes. However, existing methods neglect the importance of depth information in large-scale scene reconstruction and also fail to produce accurate depth. (b) inaccurate pose estimation Current methods struggle with pose estimation due to the error accumulation in large-scale scenes; (c) limited scene representation capability Existing approaches using a single, global model, which hinders scalability to larger scenes and longer sequences.
@@
To this end, we propose an incremental joint learning framework for depth, pose estimation, and scene implicit representation in large-scale scenes. These three elements are closely interconnected: depth information is crucial for geometric representation of the scene. It can guide the neural point sampling and enhance scene geometry consistency. Accurate pose estimation is also vital for scene representation. So, we design a joint learning framework for these three elements. The depth estimation network adopts a vision transformer-based network to encode input images into high-level features. The transformer backbone utilizes a global receptive field to process features at every stage, which provides a more precise estimation. For the pose estimation network, a feature-metric bundle adjustment (FBA) method is proposed for robust pose estimation. Compared to the existing BA methods, our FBA method effectively avoids the sensitivity of pixel-wise noise variation, enabling robust localization in large-scale scenes. For scene implicit representation, existing methods usually use an MLP to encode the scene, leading to poor scene representation capability. We propose an incremental scene representation method. We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation. In local radiance fields, unlike the coordinate-based MLPs representation, we incorporate triplane with three corresponding axes XYZ. Our feature plane-based scene representation method effectively improves both the accuracy and efficiency of scene representation. Compared to the feature grid approach, our method substantially reduces the space complexity from cubic ($\mathcal{O}(n^{3})$) to square ($\mathcal{O}(n)$). The incremental scene representation method can significantly enhance the scene representation capability in large-scale scenes.
@@
  • A novel incremental joint learning framework is proposed for accurate depth estimation, pose estimation, and large-scale scene reconstruction.
  • We design an feature-metric bundle adjustment method for the pose estimation module. Our method extracts multi-level features with pixel-wise confidences and incorporates a coarse-to-fine BA strategy to achieve robust and accurate pose estimation in large-scale scenes.
  • An incremental scene representation method is designed, which dynamically initiates local radiance fields trained with frames within a local window. We also incorporate triplane based scene representation with three corresponding axes for accurate and efficient scene reconstruction. This enables accurate scene reconstruction in arbitrarily long video and large-scale scenes.
@@ II Related work
Depth estimation [9], pose estimation [10, 11, 12], and scene representation based on monocular images have been a major problem in robotic perception and navigation [13]. Here, we summarize some relevant works.
@@ II-A Joint Learning of Depth and Pose
Depth and pose are closely interconnected components. Some early works [14] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images. [15] use synthetic images to help the disparity learning. Recently, some works have focused on self-supervised learning to jointly estimate pose and depth.  [16, 17, 18, 19] and [20] jointly optimize pose and depth via photometric reprojection error. [21] focus on the scale ambiguity for monocular depth estimation with normalization. However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency.
@@ II-B Implicit Scene Representation
With the proposal of NeRF [1], many researchers explore combining this implicit scene representation into various applications, such as autonomous driving [22, 23, 24, 25], and robot navigation. [26] integrates neural radiance fields with navigation networks end-to-end as a memory structure for the first time, highlighting the potential applications of neural radiance fields in downstream control domain tasks. [27] design a waypoint-based task decomposition strategy which effectively enhances the reliability of learning-based navigation. The key idea of NeRF is to model the radiance field of a scene with a multi-layer perceptron (MLP). However, the representation of volume densities can not commit the geometric consistency, leading to poor surface prediction for reconstruction tasks. UNISURF [28] and NeuS [29] are proposed to combine world-centric 3D geometry representation with neural radiance fields. They replace the volume density with Signed Distance Field (SDF) values, which can further improve the surface reconstruction.
@@ II-C NeRF with Pose Optimization
The original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments. They require pre-computed camera parameters obtained from SFM and SLAM algorithms [30, 31, 32, 33, 34, 35]. Recently, many researchers have focused on implicit scene representation without pose prior. iNeRF [36] is the first work focused on pose estimation and scene representation. They estimate the camera pose for novel view images with a pre-trained NeRF model. NeRFmm [37] jointly optimised as learnable parameters with NeRF training, through a photometric reconstruction. BARF [2] proposes a coarse-to-fine Bundle Adjustment method to estimate camera poses with neural 3D representation. [38, 39, 40, 41, 42] combine semantic segementation methods [43, 44] with NeRF to improve the scene representation and camera tracking. Some other methods [45, 46, 47, 48] use multiple submaps to represent the scene. NoPe-NeRF [7] and LocalRF [8] are the most relevant work to ours. NoPe-NeRF learns scale and shift parameters during training and incorporates undistorted monocular depth prior. However, they often fail in large-scale scene representation due to the inaccurate depth estimation, the accumulated pose drift, and the limited scene representation capacity. Their pose estimation module are not robust and accurate in large-scale scenes, resulting in increased pose drifts.
@@ III Joint learning of depth, pose, and implicit scene representation
In this work, we propose a joint learning framework for depth, pose estimation, and implicit scene representation. The architecture of our framework is shown in Fig. 2. The system only takes in monocular RGB images $I\in H\times W\times 3$. There are three networks in our system: depth estimation network, pose estimation network, and scene representation network. We will introduce our three networks separately.
@@ III-A Depth Estimation
In our depth network, we choose the vision transformer network as our backbone. Vision transformer network [49] has shown strong feature learning capabilities. Most large vision models, such as Segment Anything Model [50], choose it as their backbone. Inspired by this, we leverage vision transformers in place of convolutional neural networks for dense depth prediction. We maintain an encoder-decoder structure for our depth network. For the input images $I\in H\times W\times 3$, these images are decomposed into a sequence of 2D patches, where $(H,W)$ is the height and width of the original image, $(P,P)$ is the resolution of each image patches. We use ResNet50 with group norm and weight standardization to encode these image patches. Our vision transformer module transforms these patches using sequential blocks of multi-headed self-attention (MSA).
@@ III-B Feature-metric BA Pose Estimation
In terms of pose estimation, we propose a Feature-metric bundle adjustment network. We show the architecture of our network in Fig. 2. Inspired by [51], each image is encoded into high-level features and uses FBA to get an accurate and robust pose estimation. For every consecutive frames $I_{a},I_{b}$, we use a Unet-like encoder-decoder network to extract $D_{l}$ dimensional feature map $\mathbf{F}^{l}\in\mathbb{R}^{W_{l}\times H_{l}\times D_{l}}$ at each level $l\in\{L,\ldots,1\}$. Those multi-channel image features with decreasing resolution can encode richer semantic information and larger spatial information of the image. Our network can also simultaneously estimate the confidence of each pixel to select the robust features for pose estimation. Compared with the original BA method, this learned representation is robust to large illumination or viewpoint changes, which can provide accurate pose estimation in indoor and outdoor environments. Despite poor initial pose estimates, it can also provide meaningful gradients for successful estimation. Those joint learning methods with classical direct alignment [7, 8] operate on the original RGB image, which is not robust to large-scale scenes and long-term changes, and resorts to Gaussian image pyramids, which still largely limits the convergence to frame-to-frame tracking.
@@
where $\rho(\cdot)$ is a robust cost function with derivative $\rho^{\prime}$, and $\omega_{a}^{i}$ is a per-residual weight. We use the Levenberg-Marquardt (LM) algorithm to iteratively optimize this nonlinear least-squares cost from an initial estimate. Follow [51], each pose update is parametrized on the SE(3) manifold using its Lie algebra.
@@
where $\lambda$ is the damping factor and H is the Hessian matrices. Then, the final pose can be defined by left-multiplication on the manifold SE(3).
@@
2:  $p\leftarrow 1$; /$\ast$/ ∗ First camera pose index, P denotes the length of sequence/$\ast$/∗ /
@@
7:     $m\leftarrow 0$
8:     while $m<100$ do
9:        Optimize $(\theta_{j})$; /$\ast$/ ∗ Optimize the first local radiance field with the first frame/$\ast$/∗ /
10:        $m\leftarrow m+1$;
11:     end
12:     while $\{R,t\}_{j}<Bound_{j}$ do
13:        $q\leftarrow q+1$;
14:        $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; /$\ast$/ ∗ Append a pair of pose R,T at the end of the trajectory $\ast$/∗ /
@@
16:     end
17:     if $q<P$ then
18:        $j\leftarrow j+1$; /$\ast$/ ∗ A new local radiance field index/$\ast$/∗ /
19:        $\theta_{j}\leftarrow InitializeLR()$ /$\ast$/ ∗ Initialize a new local radiance field/$\ast$/∗ /
20:        $t_{j}\leftarrow t_{q}$; /$\ast$/ ∗ Centered around the last pose/$\ast$/∗ /
21:        $p\leftarrow p+n$; /$\ast$/ ∗ Stop considering the frames in last local radiance field/$\ast$/∗ /
22:     end if
23:     Repeat until all frames are registered.
24:  end
@@
where $\gamma$ is the Huber cost, L denotes the different feature level. With coarse-to-fine pose estimation, low-resolution feature maps are responsible for the robustness of the pose estimation, while finer features enhance its accuracy. This feature-metric BA loss weights the supervision of the rotation and translation adaptively for each training example. It greatly enhances the accuracy and robustness of our pose network in large-scale scenes.
Figure 3: We illustrate the differences between our sampling method and original sampling method. Our method makes better use of the estimated depth information to guide the ray sampling process.